OpenStudy (anonymous):

Chris is saving money to buy a game. So far he has saved $25, which is three-fourths of the total cost of the game. How much does the game cost?

OpenStudy (anonymous):

@kelliegirl33

OpenStudy (anonymous):

And Rewrite the expression without using a negative exponent. Simplify your answer as much as possible. 6y^-4

OpenStudy (anonymous):

hello? @agent0smith

OpenStudy (anonymous):

@.Sam.

sam (.sam.):

Think of it like this, the \(\large\frac{3}{4}\) of the total price "X", is $25. What is the total price, X? So, you could setup the equation like this \[\frac{3}{4} \times X=25 \] Then what is X? \[X=?\]

OpenStudy (anonymous):

24.25 :/

sam (.sam.):

Nope

OpenStudy (anonymous):

the only other thing im getting is 33.33

sam (.sam.):

Yup, correct

OpenStudy (anonymous):

So, that answer for that is 33.33?

sam (.sam.):

In dollars then yes

OpenStudy (anonymous):

wait wait wait. i looked at it and i typed 25 instead of 12. so now that makes it $16

OpenStudy (anonymous):

?*

sam (.sam.):

Hm?

OpenStudy (anonymous):

I wrote a 25 instead of a 12. Does that now make the answer 16

sam (.sam.):

If $12 is correct, not $25, then yeah, you're correct